OpenStudy (anonymous):

solve for x (6x-10)+(x-5)+(x-5)=180

OpenStudy (malcolmmcswain):

6x-10+x-5+(x-5)=180 Refine. 8x-20=180 Add 20 to both sides. 8x-20+20=180+20 8x=200 Divide both sides by 8. 8x/8=200/8 x=?